Calculate ∆Gº for the reaction, CH4(g)+2O2(g)→CO2(g)+2H2O(g), where ∆Gfº=-50.8 kJ/mol for CH4(g), -394 kJ/mol for CO2(g), and -229 kJ/mol for H2O(g):

572 kJ

-801 kJ

-572 kJ

801 kJ
Answer #1

DHwan = EDH formation of products - EDH formation of reactants A Gryn = {AG formation of Products - E AG formation of reactans hryn = DG+ (00g) +20Gf (H₂o) - Alf (CH4) - 2D G tla) - A Grxu = -394 +21-2290 - (-508)-210) A Gryn = -394 -458 +50.8 Ahox =